OpenStudy (anonymous):

identify the vertex and the y-intercept of the graph of the function y=(x-3)^2-5. Please explain how to do this!!

OpenStudy (kainui):

Are you in calculus or algebra?

OpenStudy (anonymous):

Algebra.

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

for the y intercept set x to 0. y= (0-3)^2 -5. Which will equal 4. So y=4

OpenStudy (anonymous):

So whenever i am finding y-intercept i make x 0?

OpenStudy (anonymous):

what about the vertex?

OpenStudy (anonymous):

Yes give me a second on the other.

OpenStudy (anonymous):

okay thanks a lot

OpenStudy (anonymous):

would the vertex be (3,5)

OpenStudy (anonymous):

The vertex is a maximum point on a graph. This graph is going to be an upward parabola.

OpenStudy (anonymous):

Is that the graph you plotted

OpenStudy (anonymous):

i didnt make a graph

OpenStudy (anonymous):

I do believe it is (3,-5). Would you care to share how you did it

OpenStudy (anonymous):

I did this y = a(x - h)^2 + k, where (h, k) is the vertex.